# Trial Length, Pricing, and Rationally Inattentive Customers

### v2

@@ 5 Shannon-Cost Specialization and Main Predictions
To derive sharp and empirically testable predictions, we now leverage the specific functional forms from our model: the Shannon information cost, which yields the logistic choice probability for $q^{*}$, and the hyperbolic decay function for attention sensitivity, $\lambda(T)$. Crucially, we also invoke our refinement assumptions, Assumptions 1 and 2, to ensure the model is tractable and well-behaved. This specialization allows us to move from the general qualitative results of the previous section to specific, quantitative predictions about the firm’s optimal contract and its response to policy shocks.
@@ 5.1 Optimal Renewal Price at a Fixed Trial Length
Lemma 2 (Full First-Order Condition for Price).
@@
$0=\;$ $\underbrace{\bigl{[}1-F(P)-Pf(P)\bigr{]}}_{\text{Standard Marginal Profit}}+\underbrace{\bigl{[}1-q^{*}\bigr{]}\bigl{\{}F(P)+Pf(P)\bigr{\}}-PF(P)\lambda\,q^{*}(1-q^{*})}_{\text{Marginal Inattentive Profit}},$ (FOC-Price)
with $\lambda=\lambda(T)$ and $q^{*}=q^{*}(P,\lambda)$.
This first-order condition precisely characterizes the firm’s complex pricing trade-off. The first bracket, $[1-F(P)-Pf(P)]$, is the standard marginal profit from happy subscribers, which is negative in the relevant region where the firm operates. It captures the classic tension between the revenue gained from a higher price and the revenue lost from consumers who no longer purchase. The second part of the equation represents the marginal profit from the inattentive segment, which is the novel contribution of our model. It has three distinct components: (1) the gain from adding a marginal consumer to the inattentive pool (the $F(P)$ term), (2) the gain from charging a higher price to the existing inattentive pool (the $Pf(P)$ term), and (3) the loss from the fact that a higher price incentivizes more consumers to pay attention and successfully cancel (the final term containing $-\lambda q^{*}(1-q^{*})$). The optimal price $P^{*}$ balances these competing effects. Unlike simpler models, this FOC explicitly shows how the optimal price depends not just on the demand curve, but also on the cognitive environment ($\lambda$) faced by consumers.
@@
The IFR property imposed by Assumption 2 is a standard regularity condition in industrial organization and marketing that ensures the firm’s profit function is single-peaked with respect to price (Lariviere and Porteus, 2006). This is a crucial technical condition. Without it, the firm might face a profit function with multiple local maxima, making it impossible to definitively characterize the optimal price. By ensuring uniqueness, the IFR property allows us to confidently use the first-order condition to identify the global maximum and to conduct well-defined comparative statics. It guarantees that for any trial length $T$ the firm might consider, there is a single, unambiguous best-response price $P^{*}(T)$.
@@ 5.2 Trial Length and Price as a Joint Choice
$\lambda^{\text{crit}}:=\sup_{P\in(0,1)}\frac{f(P)}{1-F(P)}.$
This value captures the maximum pricing power the firm would have in a standard market without inattention, as the inverse hazard rate is a key component of the Lerner index markup. The condition that follows, $\lambda_{0}>\lambda^{\text{crit}}$, essentially means that the attention-based effects must be sufficiently strong to overcome the standard pricing incentives of the firm. If attention is not a powerful enough factor, the firm’s behavior will be dominated by traditional monopoly pricing logic.
@@
If $\beta>0$ and $\lambda_{0}>\lambda^{\text{crit}}$, then $\partial P^{*}(T)/\partial T>0$ for all interior $T$.
This proposition establishes a strategic complementarity between the two main levers of the firm’s contract. It predicts that firms offering longer trial periods will also set higher renewal prices. This finding has significant implications for both marketing strategy and market analysis. It suggests that a firm’s choice of trial length is not independent of its pricing strategy but is rather an integral part of a coherent plan to segment and monetize its user base.
The economic intuition is that the trial length $T$ functions as a strategic parameter that tunes the importance of the inattentive revenue channel. As the firm increases $T$, consumer attention $\lambda(T)$ decays, making the revenue from "accidental" renewals a larger and more important component of total profit. The firm’s optimal strategy then pivots to better capitalize on this channel. By raising the price $P$, the firm increases the amount of revenue extracted from each consumer who forgets to cancel. While a higher price deters some willing subscribers, this effect is dominated by the enhanced profitability of the inattentive segment when the inattention channel is strong.
@@
This proposition confirms that the fundamental trade-off identified in Section 4 leads to a well-defined optimal contract even when the price is chosen optimally. The firm jointly optimizes over both $T$ and $P$. The IFR property from Assumption 2 ensures the problem is well-behaved, guaranteeing that a unique solution exists. This provides a complete, micro-founded explanation for the contract structures we observe in the market, where trials are neither zero-length nor infinitely long.
@@ 5.3 Policy Analysis: The "Click-to-Cancel" Effect
Our framework, now on a solid mathematical footing, allows for a nuanced analysis of common policy interventions. We revisit the "click-to-cancel" regulation, which mandates that online cancellation must be as simple as the sign-up process. We model this as an exogenous increase in attention sensitivity, $\lambda(T)\to\gamma\lambda(T)$ for $\gamma>1$. Such a policy effectively lowers the cognitive cost for all consumers to manage their subscriptions.
@@
    Under Assumption 1 (iso-elastic tail with elasticity $\varepsilon$), the sign of the price change simplifies to: $\operatorname{sign}(dP^{*}/d\gamma)=\operatorname{sign}(1-\varepsilon).$
@@
The third result, however, reveals a crucial and potentially unintended consequence that policymakers must consider. The effect of the regulation on price depends critically on the market structure, specifically the demand elasticity of the core subscriber base. Our model predicts that if the demand from these "happy subscribers" is inelastic ($\varepsilon<1$), firms will react to the loss of inattentive revenue by raising prices. Since a wealth of empirical evidence from the marketing and IS literature suggests that the demand of an existing, loyal customer base is indeed often inelastic (Dubois et al., 2006), this is the most likely outcome in many real-world markets.
@@ 6 Extension: Paid Trials and the Dilution Principle
We now extend the baseline model to consider a more general contract that includes a non-zero introductory price, $P_{0}>0$. This is a common strategy in many markets, seen in offers like "\$1 for the first three months" for streaming services or discounted first-year fees for credit cards. The introduction of an upfront fee fundamentally changes the firm’s problem by adding a screening dimension to the contract, which interacts in a non-trivial way with the attention-exploitation mechanism. The contract is now a triplet $C=(T,P_{0},P)$, and the firm must jointly optimize all three components. This extension allows us to understand the strategic choice between offering a "free trial" versus a "paid trial."
@@ 6.1 Sign-up Technology
[…]
@@
Assumption 3 (A2 , Iso-elastic Sign-up Demand).
For some results requiring a specific functional form, we will invoke the assumption that sign-up demand is iso-elastic, such that $\eta(P_{0})=\alpha P_{0}^{-\theta}$ with constants $\alpha>0$ and $\theta>0$. Under this assumption, the elasticity is constant: $\varepsilon_{0}(P_{0})\equiv\theta$.
@@ 6.3 First-Order Conditions and Strategic Interaction
Under the iso-elastic assumption (A2), this simplifies to the constant-elasticity markup rule:
@@ 7.1 Managerial Implications
For managers and marketing strategists, our model highlights that contract design is a multi-faceted problem where pricing, trial duration, and consumer psychology are deeply intertwined.
7.1.1 Contract Design as a Coherent Strategic Choice
The most critical managerial takeaway is that contract elements must be optimized jointly, not in isolation. Our finding that trial length ($T$) and renewal price ($P$) are strategic complements (Proposition 7) provides clear guidance. A strategy built around a long trial period, designed to leverage consumer inertia, should be logically paired with a relatively high renewal price. This contract structure is tailored to maximize revenue from the inattentive segment of the market. It is most effective for products with mass appeal that can attract a wide variety of consumers into the trial funnel.
@@ 7.1.1 Contract Design as a Coherent Strategic Choice
7.1.2 The Hidden Cost of Consumer Annoyance
Firms often view free trials solely through the lens of customer acquisition cost. Our model introduces a more subtle, but equally important, cost: the IR-Slack (Definition 3). This represents the erosion of consumer utility, and thus willingness to sign up, that occurs when a contract is perceived as being cognitively burdensome. A 90-day trial may seem more generous than a 30-day trial, but savvy consumers may correctly perceive it as a commitment that is harder to manage and more likely to result in an unwanted charge. This "annoyance cost" can depress sign-up rates, a phenomenon empirically verified by Miller et al. (2023). The optimal trial length $T^{*}$ (Proposition 8) is precisely the point that balances the marginal revenue from increased forgetfulness against this marginal loss in offer attractiveness. Managers should therefore be cautious about extending trial periods excessively, as they may win inattentive revenue at the expense of scaring away a valuable segment of organized, forward-looking consumers.
7.1.3 Choosing the Right Acquisition Funnel: Free vs. Paid Trials
The extension to paid trials (Section 6) offers a clear framework for deciding between a free and a low-cost introductory offer. The choice depends critically on the firm’s strategic objective.
@@ 7.1.2 The Hidden Cost of Consumer Annoyance
Strategy 2: Screen for High-Quality Customers. A paid trial ($P_{0}>0$) is the superior choice. Even a nominal fee can act as a powerful screening mechanism, filtering out less committed users and selecting for those with a higher intrinsic valuation for the product. This leads to a subscriber base with lower churn and higher lifetime value. The Dilution Principle (Proposition 12) advises that this strategy should be paired with a shorter trial period, as the need to exploit inattention is reduced. This is more common for B2B software or specialized financial news services.
@@ 7.2 Policy Implications
7.2.1 A Formal Framework for Analyzing "Dark Patterns"
Regulators are increasingly concerned with "dark patterns," which are user interface and contract designs that nudge consumers into choices they would not otherwise make (Luguri and Strahilevitz, 2021). A long automatic-renewal trial paired with an opaque and cumbersome cancellation process is a prime example of a contractual dark pattern. Our model formalizes the economic incentive for such a design: it profits by lowering the consumer’s effective attention sensitivity ($\lambda$). The harm is not merely that some consumers are "tricked," but that market competition can be distorted. Firms may be incentivized to invest in obfuscation and the exploitation of cognitive biases rather than in improving product quality or lowering prices.
7.2.2 The Double-Edged Sword of "Click-to-Cancel" Mandates
Policies such as California’s Automatic Renewal Law, which mandates that consumers must be able to cancel a service online with the same ease they signed up, are a direct intervention in this market. In our model, this is an exogenous shock that increases $\lambda$ for all consumers. Our analysis in Proposition 9 yields two crucial insights for policymakers:
Intended Consequence: The policy is effective in achieving its primary goal. By making cancellation easier, it weakens the inattention channel, compelling firms to shorten trial lengths and reduce their reliance on consumer forgetfulness. This is a clear benefit for consumer welfare, as it forces firms to compete more on the intrinsic value of their service rather than on engineered friction.
@@ 7.2.2 The Double-Edged Sword of "Click-to-Cancel" Mandates
7.2.3 Mandatory Reminders and the Structure of Trials
@@
The strategic implications of such a policy would be profound. With the inattention channel completely shut down, the firm’s primary incentive for offering a long trial would vanish. The trial would become a pure customer acquisition cost. Consequently, we would expect to see trial lengths shrink dramatically, perhaps to the minimum time required for value discovery, or even to zero. While this would offer maximum protection to forgetful consumers, it might also lead to the disappearance of the generous, long-term trial offers that some consumers value for unpressured evaluation of a product. This reveals a potential tension between policies of consumer protection and the market provision of valuable promotional tools.
7.2.4 The Case for Targeted Interventions
Our result on the convexity of the inattentive loss function (Proposition 5) suggests that the social welfare gains from improving attention are greatest at the lower end of the attention distribution. In other words, helping a very disorganized consumer become moderately organized is far more impactful than helping an already organized consumer become even more so. This provides a strong argument for targeted rather than universal interventions. While broad policies like “click-to-cancel” help all consumers, their marginal benefit is highest for the most vulnerable. This could justify complementary policies aimed specifically at these groups, such as providing financial literacy education or promoting digital tools that help manage subscriptions, as these could be the most cost-effective ways to reduce the aggregate deadweight loss from consumer inattention.

### v3

@@ 5 Shannon-Cost Specialization and Main Predictions
To derive sharp and empirically testable predictions, we now leverage the specific functional forms from our model: the Shannon information cost, which yields the logistic choice probability for $q^{*}$, and the hyperbolic decay function for attention sensitivity, $\tau(T)$. Crucially, we also invoke our refinement assumptions, Assumptions˜1 and 2, to ensure the model is tractable and well-behaved. This specialization allows us to move from the general qualitative results of the previous section to specific, quantitative predictions about the firm’s optimal contract and its response to policy shocks.
@@ 5.1 Optimal Renewal Price at a Fixed Trial Length
Lemma 2 (First-Order Condition for Price).
@@
$0=\;$ $\underbrace{\bigl[1-F(P)-Pf(P)\bigr]}_{\text{Standard Marginal Profit}}+\underbrace{\bigl[1-q^{*}\bigr]\bigl\{F(P)+Pf(P)\bigr\}-PF(P)\tau\,q^{*}(1-q^{*})}_{\text{Marginal Inattentive Profit}},$ (FOC-Price)
with $\tau=\tau(T)$ and $q^{*}=q^{*}(P,\tau)$.
This first-order condition precisely characterizes the firm’s complex pricing trade-off. The first bracket, [1F(P)Pf(P)][1-F(P)-Pf(P)], is the standard marginal profit from happy subscribers, which is negative in the relevant region where the firm operates. It captures the classic tension between the revenue gained from a higher price and the revenue lost from consumers who no longer purchase. The second part of the equation represents the marginal profit from the inattentive segment, which is the novel contribution of our model. It has three distinct components: (1) the gain from adding a marginal consumer to the inattentive pool (the $F(P)$ term), (2) the gain from charging a higher price to the existing inattentive pool (the $Pf(P)$ term), and (3) the loss from the fact that a higher price incentivizes more consumers to pay attention and successfully cancel (the final term containing $\tau q^{*}(1-q^{*})$). The optimal price $P^{*}$ balances these competing effects. Unlike simpler models, this FOC explicitly shows how the optimal price depends not just on the demand curve, but also on the cognitive environment ($\tau$) faced by consumers.
@@
The IFR property imposed by Assumption 2 is a standard regularity condition in industrial organization and marketing that ensures the firm’s profit function is single-peaked with respect to price (Lariviere and Porteus, 2001). This is a crucial technical condition. Without it, the firm might face a profit function with multiple local maxima, making it impossible to definitively characterize the optimal price. By ensuring uniqueness, the IFR property allows us to confidently use the first-order condition to identify the global maximum and to conduct well-defined comparative statics. It guarantees that for any trial length $T$ the firm might consider, there is a single, unambiguous best-response price $P^{*}(T)$.
@@ 5.2 Trial Length and Price as a Joint Choice
$\tau^{\text{crit}}:=\sup_{P\in(0,1)}\frac{f(P)}{1-F(P)}.$
This value captures the maximum pricing power the firm would have in a standard market without inattention, as the inverse hazard rate is a key component of the Lerner index markup. The condition that follows, $\tau_{0}>\tau^{\text{crit}}$, essentially means that the attention-based effects must be sufficiently strong to overcome the standard pricing incentives of the firm. If attention is not a powerful enough factor, the firm’s behavior will be dominated by traditional monopoly pricing logic.
@@
If $\beta>0$ and $\tau_{0}>\tau^{\text{crit}}$, then $\partial P^{*}(T)/\partial T>0$ for all interior $T$.
The above proposition describes a strategic complementarity between the two main components of the firm’s contract. It predicts that firms offering longer trial periods will also set higher renewal prices. This finding has significant implications for both business strategy and market analysis. It suggests that a firm’s choice of trial length is not independent of its pricing strategy but is rather an integral part of a coherent plan to segment and monetize its user base.
In Figure 4, our simulation confirms the findings so far. From the simulation result, the Lagrangian $\mathcal{L}$ exhibits an inverted U-shape with $T$, and the optimal trial length $T^{*}$ is interior, as predicted in Proposition 3. Additionally, $P$ and $T$ are complimentary, with $T^{*}$ increasing in $P$ and vice-versa, as discussed in Proposition 7 above. [^8] The set of parameters we used for the simulation is $\{\varepsilon=0.3,\beta=0.2,\tau_{0}=1\}$. In this realistic simulation scenario, the elasticity is 30%, in line with most real applications, customers start with a reasonable attention level (e.g, approximately $100\%$ pay attention to cancellation at the moment they start the trial, at $P=8$), which gradually decays over time.
Figure 4: The Firm’s Optimal Trial Length Trade-off, Conditional on Price (P)
The economic intuition is that the trial length $T$ functions as a strategic parameter that tunes the importance of the inattentive revenue channel. As the firm increases $T$, consumer attention $\tau(T)$ decays, making the revenue from "accidental" renewals a larger and more important component of total profit. The firm’s optimal strategy then pivots to better capitalize on this channel. By raising the price $P$, the firm increases the amount of revenue extracted from each consumer who forgets to cancel. While a higher price deters some willing subscribers, this effect is dominated by the enhanced profitability of the inattentive segment when the inattention channel is strong.
@@
This proposition confirms that the central trade-off identified in Section˜4 leads to a well-defined optimal contract even when the price is chosen optimally. The firm jointly optimizes over both $T$ and $P$. The IFR property from Assumption 2 ensures the problem is well-behaved, guaranteeing that a unique solution exists. This provides a complete, micro-founded explanation for the contract structures we observe in the market, where trials are neither zero-length nor infinitely long.
@@ 5.3 Policy Analysis: The "Click-to-Cancel" Effect
Our framework, as established in the previous sections, allows for a nuanced analysis of common policy interventions. In this section, we apply it to the the "click-to-cancel" regulation, recently proposed by the FTC as well as California state legislative, which mandates that online cancellation must be as simple as the sign-up process. We model this as an exogenous increase in attention sensitivity, $\tau(T)\to\gamma\tau(T)$ for $\gamma>1$. Such a policy effectively lowers the cognitive cost for all consumers to manage their subscriptions.
@@
    Under Assumption 1 (Isoelastic tail with elasticity $\varepsilon$), the sign of the price change simplifies to: $\operatorname{sign}(dP^{*}/d\gamma)=\operatorname{sign}(1-\varepsilon).$
@@
The third result, however, reveals a crucial and potentially unintended consequence that policymakers must consider. The effect of the regulation on price depends critically on the market structure, specifically the demand elasticity of the core subscriber base. Our model predicts that if the demand from these "happy subscribers" is inelastic ($\varepsilon<1$), firms will react to the loss of inattentive revenue by raising prices. Since a wealth of empirical evidence from extant literature suggests that the demand of an existing, loyal customer base is indeed often inelastic (Krishnamurthi and Raj, 1991), this is the most likely outcome in many real-world markets.
@@ 6 Extension: Paid Trials and the Dilution Principle
We now extend the baseline model to consider a more general contract that includes a non-zero introductory price, $P_{0}>0$. This is a common strategy in many markets, seen in offers like "\$1 for the first three months" for news subscription services (e.g., the New York Times, The Washington Post) or discounted first-year fees for credit cards. The introduction of an upfront fee fundamentally changes the firm’s problem by adding a screening dimension to the contract, which interacts in a non-trivial way with the attention-exploitation mechanism. The contract is now a triplet $C=(T,P_{0},P)$, and the firm must jointly optimize all three components. This extension allows us to understand the strategic choice between offering a "free trial" and a "paid trial."
@@ 6.1 Sign-up Technology
$\varepsilon_{0}(P_{0}):=-\eta^{\prime}(P_{0})\,\frac{P_{0}}{\eta(P_{0})}>0\quad\text{(elasticity).}$
@@
Assumption 3 (A3: Isoelastic Sign-up Demand).
For some results requiring a specific functional form, we will invoke the assumption that sign-up demand is isoelastic, such that $\eta(P_{0})=\alpha P_{0}^{-\theta}$ with constants $\alpha>0$ and $\theta>0$. Under this assumption, the elasticity is constant: $\varepsilon_{0}(P_{0})\equiv\theta$.
@@ 6.3 First-Order Conditions and Strategic Interaction
Under the isoelastic assumption (A2), this simplifies to the constant-elasticity markup rule:
@@ 7.1 Managerial Implications
For managers, our model highlights that contract design is a multi-faceted problem where pricing, trial duration, and consumer psychology are deeply intertwined.
Contract Design as a Coherent Strategic Choice
The most critical managerial takeaway is that contract elements must be optimized jointly, not in isolation. Our finding that trial length ($T$) and renewal price ($P$) are strategic complements (Proposition˜7) provides clear guidance. A strategy built around a long trial period, designed to exploit consumer inertia, should be logically paired with a relatively high renewal price. This contract structure is tailored to maximize revenue from the inattentive segment of the market. It is most effective for products with mass appeal that can attract a wide variety of consumers into the trial funnel.
@@
The Hidden Cost of Consumer Annoyance
Firms often view free trials solely through the lens of customer acquisition cost. Our model introduces a more subtle, but equally important, cost: the IR-Slack (Definition˜2). This represents the erosion of consumer utility, and thus willingness to sign up, that occurs when a contract is perceived as being cognitively burdensome. A 90-day trial may seem more generous than a 30-day trial, but savvy consumers may correctly perceive it as a commitment that is harder to manage and more likely to result in an unwanted charge. This "annoyance cost" can depress sign-up rates, a phenomenon empirically verified by Miller et al. (2023). The optimal trial length $T^{*}$ (Proposition˜8) is precisely the point that balances the marginal revenue from increased forgetfulness against this marginal loss in offer attractiveness. Managers should therefore be cautious about extending trial periods excessively, as they may win inattentive revenue at the expense of scaring away a valuable segment of organized, forward-looking consumers.
Choosing the Right Acquisition Funnel: Free vs. Paid Trials
The extension to paid trials (Section˜6) offers a framework for deciding between a free and a low-cost introductory offer. The choice depends critically on the firm’s strategic objective.
@@
Strategy 2: Screen for High-Quality Customers. A paid trial ($P_{0}>0$) is the superior choice. Even a nominal fee can act as a powerful screening mechanism, filtering out less committed users and selecting for those with a higher intrinsic valuation for the product. This leads to a subscriber base with lower churn and higher lifetime value. The Dilution Principle (Proposition˜12) advises that this strategy should be paired with a shorter trial period, as the need to exploit inattention is reduced. This is more common for B2B software or specialized financial news services.
@@ 7.2 Policy Implications
A Formal Framework for Analyzing "Dark Patterns"
Regulators are increasingly concerned with "dark patterns," which are user interface and contract designs that nudge consumers into choices they would not otherwise make (Luguri and Strahilevitz, 2021). A long automatic-renewal trial paired with an opaque and cumbersome cancellation process is a prime example of a contractual dark pattern. Our model formalizes the economic incentive for such a design: it profits by lowering the consumer’s effective attention sensitivity ($\tau$). The harm is not merely that some consumers are "tricked," but that market competition can be distorted. Firms may be incentivized to invest in obfuscation and the exploitation of cognitive biases rather than in improving product quality or lowering prices.
The Double-Edged Sword of "Click-to-Cancel" Mandates
Policies such as California’s Automatic Renewal Law, which mandates that consumers must be able to cancel a service online with the same ease they signed up, are a direct intervention in this market. In our model, this is an exogenous shock that increases $\tau$ for all consumers. Our analysis in Proposition˜9 yields two crucial insights for policy makers:
Intended Consequence: The policy is effective in achieving its primary goal. By making cancellation easier, it weakens the inattention channel, compelling firms to shorten trial lengths and reduce their reliance on consumer forgetfulness. This is a straightforward benefit for consumer welfare, as it forces firms to compete more on the intrinsic value of their service than on engineered friction.
@@ A Formal Framework for Analyzing "Dark Patterns"
Mandatory Reminders and the Structure of Trials
@@
The strategic implications of such a policy would be profound. With the inattention channel completely shut down, the firm’s primary incentive for offering a long trial would vanish. The trial would become a pure customer acquisition cost. Consequently, we would expect to see trial lengths shrink dramatically, perhaps to the minimum time required for value discovery, or even to zero. Although this would offer maximum protection to forgetful consumers, it might also lead to the disappearance of the generous long-term trial offers that some consumers value for an low-pressure evaluation of a product, lowering consumer learning. This reveals a potential tension between consumer protection regulations and the market provision of valuable promotional tools.
The Case for Targeted Interventions
Our result on the convexity of the inattentive loss function (Proposition˜5) suggests that the social welfare gains from improving attention are greatest at the lower end of the attention distribution. In other words, helping a very disorganized consumer become moderately organized is far more impactful than helping an already organized consumer become even more so. This provides a strong argument for targeted rather than universal interventions. While broad policies like “click-to-cancel” help all consumers, their marginal benefit is highest for the most vulnerable. This could justify complementary policies aimed specifically at these groups, such as providing financial literacy education or promoting digital tools that help manage subscriptions, as these could be the most cost-effective ways to reduce the aggregate deadweight loss from consumer inattention.